OpenStudy (anonymous):

A student on a sled travels down a frictionless incline and arrives at the bottom with a velocity of va. If 5 additional students jump on the sled at the top of the incline and repeat the trip a) what will be the new velocity of the sled when it arrives aat the bottom, relative to va b) support your answer to part a mathematically. show all steps

OpenStudy (anonymous):

velocity will be the same. Potential energy is being converted into kinetic energy, both depend equally on the mass. Ep = mgh and Ek = 1/2mV^2 so gh = 1/2 V^2 or V = (2.gh)^0.5